Hot exhaust gases are used in a shell-and-tube exchanger (1 shell and 20 tube passes) to heat 2.5 kg/s of water from 35 to 84Please answer clearly and neatly so I can read your work. I will give a good rating! Thank you (:
Answer #1

flow Heat exchange solution Let us assume parallel Thi 200 Aia. mm ho wc W I r 10=842 To =35c water mc2-5leg! (pc=41855 / regme small = 4791 (МСЬја (очCars in c= 0.458 - CHCINTO Ex He=t II (H0-458) NTU 8-6484 = te Hou58 AG NTU=2 »NTU= UA (me small 2=